OpenStudy (anonymous):

Simplify the following product of complex numbers.(-3 + 2i) · (-4 – 8i). I got 12-16i is this correct or am I doing something wrong?

OpenStudy (anonymous):

think there is something wrong here

OpenStudy (anonymous):

ok could you please explain what I am doing wrong oh wait it is 12-16i^2

OpenStudy (anonymous):

\[-3\times -4+2\times 8=28\]

OpenStudy (anonymous):

12-24i-8i+16, 28-32i, or am I being dumb?

OpenStudy (anonymous):

omg I am so confused now

OpenStudy (anonymous):

\[(-3 + 2i) · (-4 – 8i)\] \[(-3\times -4+2\times 8)+(2\times -4-3(-8))i\]

OpenStudy (anonymous):

if that is confusing you can multiply it out just like you would \[(-3 + 2x) · (-4 – 8x)\]

OpenStudy (anonymous):

But remember x^2, i^2 is -1.

OpenStudy (anonymous):

aka first outer inner last \[12 +24i-8i-16i^2\] \[12+16i+16\] \[28+16i\]

OpenStudy (anonymous):

Oh right, I failed to multiply -3 and -8i correctly. Woops. Lol.

OpenStudy (anonymous):

I multiplied -3*-4=12 then got multiplied 2i *8i=16i^2 I am completely conpletely confused feel like an idiot math has never had me so dumb founded

OpenStudy (anonymous):

I get it!!!! I was not solving the whole problem got it THANK YOU THANKYOU SO MUCH!!!!